Consider the test of H0 : σ2-5 against H1 : σ2 < 5. Approximate the P-value for the following test statistic. 215.2 and n 12 0.01 < P-value < 0.05 0.25< P-value 0.75 0.5< P-value < 0.9 0.1 < p-value < 0.5 O 0.05<P-value< 0.09
Answer #1

Solution :

Given that,

This is the left tailed test .

\chi2 = 15.2

n = 12

df = n - 1 = 12 - 1 = 11

Using chi square table the p-value is ,

0.5 < p-value < 0.9